### v1

# IPPP/25/40

Abstract
The primordial power spectrum of matter density perturbations contains highly valuable information about new fundamental physics, in particular cosmological inflation, but is only very weakly constrained observationally for small cosmological scales $k\gtrsim 3\,{\rm Mpc}^{-1}$. We derive novel constraints, $\mathcal{P}_{\mathcal{R}}(k)\lesssim 5\cdot 10^{-6}$ over a large range of such scales, from the formation of ultracompact minihalos in the early universe. Unlike most existing constraints of this type, our results do not rest on the assumption that dark matter can annihilate into ordinary matter.
I Introduction
The power spectrum of primordial density fluctuations, $\mathcal{P}(k)$, is a key quantity in cosmology. It describes the initial conditions of the Universe right after the big bang, tiny perturbations in an almost homogeneous plasma that eventually evolved into all large-scale cosmological structures that we can see today. Exquisite observations of the cosmic microwave background radiation (CMB) [1] and large-scale structure, in particular through the Lyman-$\alpha$ forest [2, 3], tightly constrain $\mathcal{P}(k)$ on length scales larger than $\sim 1$ Mpc. These observations are consistent with the simplest models of cosmological inflation [4], which represents a remarkable success for our leading explanation of how those primordial perturbations were seeded initially. $\mathcal{P}(k)$ is thus not only crucial for understanding the universe at large scales, but also contains invaluable information about fundamental new physics, connected to much higher energy scales than what is accessible by earth-bound experiments.
Fully exploiting this potential would require to extend our knowledge of $\mathcal{P}(k)$ to sub-Mpc scales, but the heap of information that may be buried there remains largely inaccessible at present. At the smallest scales, in particular, the non-observation of primordial black holes (PBHs) provides our only probe of $\mathcal{P}(k)$; this leads to rather mild constraints, about seven orders of magnitude weaker than what is observed at large scales [5, 6]. Ultracompact minihalos (UCMHs) form much more frequently from rare density fluctuations in the early universe than PBHs [7, 8, 9]. Assuming that dark matter (DM) can pair-annihilate into standard model particles, like in the case of the much-discussed weakly interacting massive particle (WIMP) explanation of DM [10, 11], this translates into very stringent small-scale constraints on $\mathcal{P}(k)$ [12, 9, 13, 14, 15, 16, 17, 18]. Future pulsar timing array observations of UCMH transients might constrain $\mathcal{P}(k)$ independent of this assumption [19, 20]. Later forming DM halos may also lead to various constraints at sub-Mpc scales, subject to assumptions about the late-time evolution and properties of these objects [21, 22, 23, 24, 25, 26, 27]. Finally, spectral distortions of the CMB [28] and the induced gravitational wave background [29, 30] provide complementary constraints directly on the linear, or only mildly non-linear, power spectrum at scales down to roughly one pc.
UCMHs are only one example of extended DM objects (EDOs) that may have existed since primordial times. Recently, in fact, the possibility of more general macroscopic DM configurations, very compact yet much larger in size than PBHs, has attracted attention [31, 32, 33, 34, 35]. Even if only interacting gravitationally, EDOs would typically constitute gravitational lenses and accrete ordinary, baryonic matter; they may also disrupt other gravitational structures such as wide binaries. As it turns out, these effects lead to stringent constraints on their cosmological abundance, largely independent of specific assumptions about the EDO genesis or the microscopic nature of DM (see Ref. [36] for an up-to-date overview).
Here we combine the idea of using UCMHs to constrain the primordial power spectrum with these recent developments in the more general EDO context. We present new and significantly extended limits on EDOs, and apply them to the more specific case of UCMHs. In combination, this allows us to improve existing constraints on $\mathcal{P}(k)$ by up to an order of magnitude, over a wide range of sub-Mpc scales. Such tightened small-scale constraints will not only be decisive in further narrowing down the microphysics governing inflation (see, e.g., Ref. [37]), but also put highly relevant bounds on other mechanisms generating excess power on small scales, like dark sector phase transitions [18].
II UCMH formation and properties
Let us consider a region of comoving radius $R$ in the early universe, with an average density $\rho$ that is larger than the cosmological background density $\bar{\rho}$. For sufficiently large overdensities $\delta\equiv(\rho-\bar{\rho})/\bar{\rho}>\delta_{\chi}^{\rm min}\sim 10^{-3}$ during radiation domination, this region will collapse into virialized DM objects shortly after matter-radiation equality [7]. Since such large values of $\delta$ correspond to very rare fluctuations, the collapse typically happens in isolation and by almost radial infall, leading to the formation of highly compact objects, dubbed UCMHs. Their mass corresponds to the DM mass within the horizon $H^{-1}$ at mode entry,
$M_{\rm UCMH}\simeq\frac{4\pi}{3}\left[\rho_{\chi}H^{-3}\right]_{aH=1/R}=\frac{1}{2}M_{\rm Pl}^{2}H_{0}^{2}\Omega_{\chi}R^{3}\,,$ (1)
where $a$ is the scale factor, $M_{\rm Pl}$ the Planck mass, $\rho_{\chi}$ the DM density and we introduced $\Omega_{\chi}\equiv\rho_{\chi}/\rho_{c}=0.264$ [1], with $\rho_{c}$ being the critical density. It was originally argued that accretion would enhance $M_{\rm UCMH}$ by another factor of $a_{\rm eq}/a$ after equality, but such a growth could not consistently be reproduced in simulations [38] and we hence conservatively neglect it for most of our work.
The density profile inside these objects is very steep, $\rho_{\rm UCMH}(r)\propto r^{-\alpha}$, with $\alpha=9/4$ expected for self-similar, spherical collapse [39, 40, 41, 42]. While this value of $\alpha$ is realistic for fully isolated overdensities, numerical simulations often find less spiky profiles [43, 44, 45], even for rather early collapse [38, 46, 47]. We will hence assume $\alpha=3/2$, noting that our limits anyway only weakly depend on the exact value of $\alpha$ (see Appendix for a more detailed discussion). Unlike for the case of DM annihilation, our analysis is also insensitive to the inner cutoff radius at which $\rho_{\chi}$ plateaus because of the non-vanishing velocity distribution or DM depletion due to self-annihilation [9]. We take the forming UCMH to extend out to an effective radius where its density has dropped to three times the background density, $R(z)\simeq 21(1+z)^{-1}(M_{\rm UCMH}/M_{\odot})^{1/3}\,{\rm pc}$ [48]. The radius within which $90\%$ of the mass is enclosed is thus given by $R^{(90)}_{\rm UCMH}=(0.9)^{3/2}R(z_{c})$, with $z_{c}$ the redshift of collapse. This translates to $\rho_{\rm UCMH}(r)=\rho_{90}\big{(}r/R^{(90)}_{\rm UCMH}\big{)}^{-3/2}$ with $\rho_{90}\simeq 1.4\cdot 10^{-5}(1+z_{c})^{3}\,M_{\odot}\,{\rm pc}^{-3}$.
For Gaussian fluctuations, the cosmological abundance of UCMHs is exponentially sensitive to the critical overdensity $\delta_{\chi}^{\rm min}$ for collapse to happen. We follow Ref. [9] to calculate this value, very conservatively requiring collapse by $z_{c}=1200$. For the range of scales that will be relevant for us, $3\,{\rm Mpc}^{-1}\lesssim k\lesssim 10^{5}\,{\rm Mpc}^{-1}$, this corresponds to $1.03\cdot 10^{-2}\gtrsim\delta_{\chi}^{\rm min}\gtrsim 2.25\cdot 10^{-3}$.
III Updated EDO limits
Recently, gravitational constraints on generic EDOs have been derived from the non-observation of micro-lensing events [33, 31], gravitational wave production [34], the evaporation of wide binaries [49, 50], dynamical heating of dwarf galaxies [51], and heating of the CMB due to baryonic accretion [35]. A repository of current constraints on EDOs with different mass profiles is provided in Ref. [36]. In this work, we extend previous bounds on EDOs from accretion and introduce a new constraint from microlensing of the MACS J1149 LS1 “Icarus” star.
In previous analyses the accretion constraint was taken to be valid up to masses of $M=10^{5}M_{\odot}/\sqrt{f_{\rm DM}}$, where $f_{\rm DM}$ is the DM fraction in EDOs. Beyond this the average EDO separation is much smaller than their Bondi radius $R_{\rm B}(z)=GM/c_{s}(z)^{2}$, with $c_{s}$ being the sound speed, raising concerns about the underlying assumption of isolated objects. A priori, exceeding this limit may introduce backreaction effects (potentially underestimating constraints) or lead to overcounting the total emitted radiation (potentially overestimating them). However, this latter effect is significantly suppressed, as only the ionised matter (heated above $T_{\rm ion}=1.3\,\rm eV$ [52]) contributes to luminosity. This only occurs within the ionisation radius $R_{\rm ion}=R_{\rm B}(T_{\rm CMB}(z)/T_{\rm ion})\,\Theta(z)^{-2/3}$, where $\Theta(z)$ accounts for interactions with the background [32, 35]. Since $R_{\rm ion}\ll R_{\rm B}$ stays well below the average EDO separation across relevant redshifts, the isolated object approximation remains a conservative and valid basis for extending constraints to higher-mass EDOs.
Observations of highly magnified background stars behind galaxy clusters yield a novel independent bound on EDOs. In favourable lensing configurations, where a distant star is strongly amplified by a cluster caustic, the magnification hinges on the smoothness of the cluster’s mass profile. If DM consists of compact objects, it fragments the caustic into a network of microcaustics, modifying the lensing signal. The lensed observation of the Icarus star has so far only been used to constrain the DM fraction in PBHs [53, 54], which we explain in the Appendix. In an extended macrolens, the effective Einstein radius of a point lens at distance $D_{L}$ and Einstein angle $\theta_{E}$ is $\bar{R}_{E}=\sqrt{\mu_{t}}\,\theta_{E}\,D_{L},$ which for MACS J1149 LS1 is boosted by at least $\mu_{t}\sim 15.4$ [53]. Lenses behave as point-like when their physical size satisfies [31, 33]
$R_{\rm 90}\lesssim\bar{R}_{E}=\sqrt{\mu_{t}}\,\theta_{E}\,D_{L}\approx 2.0\times 10^{6}\bigl{(}M/M_{\odot}\bigr{)}^{1/2}R_{\odot}\,,$ (2)
using $\theta_{E}=1.8\times 10^{-6}(M/M_{\odot})^{1/2}\rm arcsec$ and $D_{L}=6.4\,\rm kpc\,arcsec^{-1}$ for this system [53]. We conservatively apply point-like microlensing bounds up to this radius, leaving a more detailed treatment to future work.
Figure 1: Present bounds on the DM fraction $f_{\rm DM}$ comprised of EDOs (blue), with fixed radii as stated in the legend, and UCMHs (red). Constraints deriving from lensing observations of the Icarus star and from CMB heating due to baryonic accretion are described in detail in the text. We further apply Galactic microlensing [31, 33] and wide binary evaporation constraints [49].
In Fig. 1 we plot the constraints on the EDO abundance discussed above (blue contours), for various fixed values of the size $R_{90}$ of these objects. We choose a density profile $\rho_{\rm EDO}\propto r^{-\alpha}$ with $\alpha=3/2$, stressing however that these constraints are not very sensitive to the precise value of $\alpha$ (see also the Appendix) and mostly depend on the underlying relation between mass and $R_{90}$ [35, 36]. In red, we overlay the constraints on objects which satisfy the UCMH mass-radius relation $R^{(90)}_{\rm UCMH}\propto M^{1/3}$. We note that UCMHs are not sufficiently compact to be probed by Galactic microlensing, see also the discussion below, and that CMB accretion constraints by far dominate at high masses.
Figure 2: Shaded regions show the presently allowed strength of a locally scale-invariant primordial curvature power spectrum $\mathcal{P}_{\mathcal{R}}(k)$, as a function of co-moving scale $k$. Our newly derived limits from UCMHs due to CMB accretion (blue), wide binary evaporation (olive) and ICARUS lensing observations (light green) are discussed in detail in the text. We also present updated constraints deriving from PTAs (black), as well as existing CMB $y$ and $\mu$ distortion constraints applied to the form of $\mathcal{P}_{\mathcal{R}}(k)$ assumed here (red). For comparison, we further show existing constraints at large scales (purple) deriving from CMB and Lyman-$\alpha$ observations [1, 55], as well as PBH bounds [5] that extend to very small scales (grey).
IV UCMH constraints on the power spectrum
In the previous section we presented updated constraints on the EDO abundance. By applying those to UCMHs, we now show how this translates to constraints on the curvature power spectrum $\mathcal{P}_{\mathcal{R}}(k)$. Assuming Gaussian fluctuations, the probability for a region of size $R$ to eventually collapse into a UCMH is given by
$\beta=\frac{1}{\sqrt{2\pi}\sigma_{\chi,H}}\int^{\delta^{\rm max}_{\chi}}_{\delta^{\rm min}_{\chi}}\exp\left[-\frac{\delta_{\chi}^{2}}{2\sigma^{2}_{\chi,H}}\right]{\rm d}\delta_{\chi}\,,$ (3)
which depends only very weakly on the upper integration boundary $\delta^{\rm max}_{\chi}\simeq\mathcal{O}(1)\gg\delta^{\rm min}_{\chi}$ corresponding to the formation of a PBH rather than an UCMH. Here, the mass variance $\sigma_{\chi,H}$ at horizon crossing of the scale $R$ is computed by convolving the power spectrum with a top-hat window function [56, 9]. For scales $k_{R}=1/R$ entering during radiation domination, it takes the form
$\sigma_{\chi,H}^{2}=\int_{0}^{\infty}\!\!{\rm d}x\,\frac{\left(\sin x-x\cos x\right)^{2}}{x^{3}}\mathcal{P}_{\mathcal{R}}(x/R)\,T_{\chi}(x/\sqrt{3})\,,$ (4)
with the transfer function $T_{\chi}$ given explicitly in Eq. (A10) of Ref. [9]. For most functional forms of $\mathcal{P}_{\mathcal{R}}(k)$, this integral is strongly dominated by $x\sim 1$, corresponding to scales $k\sim k_{R}$. For ease of comparison, we will present our constraints in terms of a power spectrum that is locally scale-invariant, $\mathcal{P}_{\mathcal{R}}(k)=\mathcal{P}_{\mathcal{R}}(k_{\mathrm{R}})$ for $k\sim k_{R}$, leading to $\sigma^{2}_{\chi,{\rm{H}}}(R)\simeq 0.91\,\mathcal{P}_{\mathcal{R}}(k_{\mathrm{R}})$. We stress, however, that Eq. (4) can, without further complications, also be used to constrain power spectra of arbitrary shapes.
Neglecting the accretion of baryonic matter, the final UCMH abundance is given as $\Omega_{\rm UCMH}=\Omega_{\rm DM}\beta$. Comparing this to the limits above leads to the constraints on $\mathcal{P}_{\mathcal{R}}(k)$ shown in Fig. 2. It is clear that CMB accretion, wide binary evaporation and the Icarus star are highly complementary, leading to competitive bounds for scales up to a factor of $\sim 10^{4}$ smaller than what can be directly measured by large-scale structure observations.
Further constraints.—
For this range of scales, it has previously been pointed out that also the lack of Compton-$y$ and $\mu$ distortions of the CMB black-body spectrum constrains scalar perturbations. We follow Ref. [57] in computing both $y$ and $\mu$ from a given power spectrum, assuming again $\mathcal{P}_{\mathcal{R}}(k)\simeq const.$ around the pivot scale $k_{R}$. For concreteness, we assume that $\mathcal{P}_{\mathcal{R}}(k)$ is nonzero only for the range $1/3\lesssim k/k_{R}\lesssim 3$. Comparing the result to the $2\sigma$ limits of $y<1.5\times 10^{-5}$ and $\mu<9\times 10^{-5}$ based on FIRAS observations [58] then leads to the constraints shown as red lines in Fig. 2. We caution, however, that $\mu$ and $y$ in general receive contributions from a significantly larger range of scales than $\sigma_{\chi,H}$ in Eq. (4), which necessarily makes any comparison with UCMH limits somewhat model-dependent (see Appendix for a more detailed discussion).
Another relevant constraint that is independent of UCMH formation derives from the fact that scalar perturbations produce a stochastic gravitational wave background (SGWB) at second order in perturbation theory [59, 60]. The computation of the resulting power spectrum in GWs, $\mathcal{P}_{h}$, involves highly oscillatory integrals. We perform those semi-analytically by adopting an averaging procedure as advocated in Refs. [61, 62], noting that the innermost integral is dominated by contributions from modes well inside the horizon, and by assuming $\mathcal{P}_{\mathcal{R}}(k)$ to be scale-invariant over the same range of scales as for CMB distortions. We then compute the resulting fractional energy density of gravitational waves today, $\Omega_{\rm GW}h^{2}(k)=(\Omega_{r}h^{2})\frac{1}{24}({k}/aH)^{2}\mathcal{P}_{h}(k,t)$, where the r.h.s. is evaluated at any time $t_{k}\lesssim t\lesssim t_{\rm eq}$ and the factor of $\Omega_{r}h^{2}=4.15\cdot 10^{-5}$ [58] accounts for the red-shifting of $\Omega_{\rm GW}$ during matter domination. We display the resulting limits on $\mathcal{P}_{\mathcal{R}}(k)$ as black lines in Fig. 2, resulting from the requirement of $\Omega_{\rm GW}h^{2}$ not exceeding the NANOGrav data [63] at 99% C.L in either of the frequency bins.
We also include a compilation of PBH bounds [5] in the figure, which remain the strongest constraints for scales $k\gtrsim 10^{6}\,{\rm Mpc}^{-1}$. At large scales, finally, the power spectrum is well measured by observations of the CMB temperature fluctuations and the Lyman-$\alpha$ forest. Here we show results from Ref. [55] for the former, based on Planck data [64], and from Ref. [3] for the latter.
V Discussion
We have demonstrated that the formation of UCMHs leads to highly competitive constraints on $\mathcal{P}_{\mathcal{R}}(k)$ down to scales $k\sim 10^{5}$ kpc. Remarkably, these constraints do not depend on interactions between DM and ordinary matter and hence apply to any collision-less and sufficiently cold DM candidate (in particular, these constraints only apply below the scale of free streaming or kinetic decoupling [65], $k<k_{D}$, for DM particles that have been in thermal equilibrium in the early universe). We stress that we have followed a very conservative approach by restricting our analysis to objects that have fully collapsed very early, before a redshift of $z_{c}=1200$. Notably, this considerably simplifies the analysis as it largely avoids complications due to potential changes in the inner density profile or tidal disruption affecting the survival probability at later formation times.
Still, even significantly smaller initial density contrasts than considered here would imply perturbations that enter the non-linear regime well before the onset of standard structure formation. While UCMHs forming at $z_{c}\lesssim 1000$ no longer efficiently contribute to the CMB accretion bounds, both wide binary evaporation and the Icarus observation therefore generally lead to significantly stronger limits than what we reported in Fig. 2. Due to the aforementioned effects, however, such more aggressive limits should be considered less robust. To a much lesser degree, our constraints are affected by other assumptions about the UCMH properties. Taking into account a potential mass growth due to baryonic accretion, in particular, would lead to somewhat stronger bounds, while assuming a shallower profile (as expected for later collapse) would result in slightly weaker constraints. For a more detailed discussion of theses aspects, we refer to the Appendix.
Before concluding, let us briefly discuss further, complementary probes of the small-scale power spectrum. To start with, microlensing is effective when the physical size of the lens is smaller than the Einstein radius $R_{E}\leq R_{E,\text{max}}=\sqrt{{GMD_{S}}/{c^{2}}}$, where $D_{S}$ is the distance to the source and $M$ the lens mass, here assumed to be an UCMH. For Galactic microlensing, this implies $M>10^{{17}}\rm M_{\odot}$, which is far above the masses to which surveys are sensitive. This explains why there are no UCMH microlensing constraints in Figs. 1 and 2 but, if a greater UCMH compactness is assumed, future microlensing observations can lead to constraints on $\mathcal{P}(k)$ [66]. In contrast, the larger source distance and effective boost in $R_{E}$ make highly magnified stars such as Icarus an effective probe. Moreover, future observations of gravitational wave lensing with both sources and lenses at cosmological distances may also lead to constraints on UCMH [67, 68]. Astrometric microlensing has also been suggested as a future probe [69], as this can be detectable even if the magnification is small.
Recently, Ref. [70] translated the Planck optical depth into limits on $\mathcal{P}_{\mathcal{R}}(k)$ at low $k\sim 3\!-\!50\ \mathrm{Mpc}^{-1}$, using a one-zone reionisation model with fixed ionising efficiency, escape fraction and IGM clumping, assuming UCMH survival until $z\sim 200$. At higher $k\sim 10\!-\!10^{3}\ \mathrm{Mpc}^{-1}$, Ref. [51] derived dynamical-heating constraints from ultrafaint dwarf galaxies complementary to the bounds shown here, however using NFW subhalos instead of the UCMHs considered here. Like the microlensing constraints, these rely on subhalo survival without significant tidal disruption or baryonic feedback.
Finally, we note that the results in this article may also have implications for isocurvature perturbations, since the formation of UCMHs is highly sensitive to non-adiabatic fluctuations on small scales. In particular, improved constraints on the small-scale power spectrum may tighten bounds on isocurvature modes that evade CMB limits but still source early structure formation.
VI Conclusions
While PBHs currently remain the only probe of the smallest cosmological scales, with $k\gtrsim 10^{-6}$ Mpc-1, there are several independent observations that put much more stringent constraints on the primordial power spectrum at slightly larger scales. Here we pointed out that very early forming UCMHs can be used to significantly tighten these bounds, in a way that is both robust and largely independent of the nature of DM. Notably, this also provides complementary bounds on PBH DM at these scales. Future optical and gravitational wave lensing observations have the potential of even further improvement. Such probes of the small-scale primordial power spectrum constitute a crucial step towards understanding the microphysics governing comic inflation, as well as more exotic processes that may generate excess power at these scales.
Acknowledgements.—
TB gratefully acknowledges support through a FRIPRO grant of the Norwegian Research Council (project ID 353561 ‘DarkTurns’). DC and SSM are supported by the STFC under Grant No. ST/T001011/1. TB and SSM are grateful to the Mainz Institute for Theoretical Physics (MITP) of the Cluster of Excellence PRISMA+ (Project ID 390831469), for its hospitality and partial support during the DMLAND2024 workshop where the idea for this work was born.
Appendix A UCMH formation
In order to be as conservative as possible, we so far only considered UCMHs that collapse very early, namely by a redshift of $z_{c}=1200$. A collapse at smaller redshifts would require smaller and thus much more likely initial density contrasts $\delta\gtrsim\delta_{c}$. On the other hand, later forming objects are less dense, with $\rho_{90}\propto(1+z_{c})^{3}$, and extend to larger radii $R_{\rm UCMH}^{(90)}\propto(1+z_{c})^{-1}$. Notably, the latter aspect does not only affect the survival probability, but even detection prospects under the (optimistic) assumption that the late-time cosmological evolution does not affect abundance and structure of these objects.
In Fig. 3 we demonstrate how this affects our limits on $\mathcal{P}_{\mathcal{R}}(k)$, noting that CMB bounds in any case require collapse before recombination. The other constraints, from wide binary evaporation and lensing of the Icarus star, clearly show the generally expected behaviour outlined above: while the constraint on the power spectrum normalization tightens as a direct consequence of the decrease in $\delta_{c}$, the mass range where the constraints apply is reduced as the physical UCMH size increases. The actual reduction of the maximally constrained value of $k$ visible in the figure is further enhanced, compare to the naively expected $\propto(1+z_{c})$, by the fact that the Icarus bound itself is mass-dependent.
In reality, collapse would happen at all the redshifts indicated in the figure, so the overall constraint on $\mathcal{P}_{\mathcal{R}}(k)$ should in principle be taken as the envelope of the curves shown here. However, we re-iterate that these constraints become less and less robust as $z_{c}$ is decreased. For late-time collapse, a more realistic computation should thus take into account both survival probability and changes in the UCMH density profile due to merger events and other effects during non-linear structure formation. This must be based on full numerical simulations and is beyond the scope of this work.
Figure 3: Dependence of the bounds on $\mathcal{P}_{\mathcal{R}}(k)$ for wide binary evaporation and Icarus constraints on the assumed redshift $z_{c}$ of collapse. CMB accretion bounds require formation before recombination.
Appendix B UCMH density profiles and growth by accretion
As stressed in the main text, UCMHs forming very early and in isolation are expected to obtain very steep density profiles $\rho\propto r^{-\alpha}$ with $\alpha=9/4$. Once either of these assumptions is not met, however, somewhat less steep profiles with $\alpha\simeq 3/2$ are more realistic. Also independently of the UCMH context it is interesting to ask how our general EDO constraints are affected by the density profile for $r<r_{90}$. We show this in Fig. 4, demonstrating that the value of $\alpha$ only has a relatively mild impact on the limit. For both the wide binary evaporation constraint and the CMB accretion constraint, this remains true as long as the objects under consideration remain at least somewhat compact, implying that such limits are dominantly driven by the relation between $M$ and $R_{90}$ [35, 36]. The microlensing efficiency for extended objects displays some sensitivity to the density profile, as was discussed in Ref. [31], where the threshold impact parameter effectively encodes the efficiency (see in particular their Fig. 3). In contrast, the conservative assumption we adopt for applying the Icarus constraint assumes no such profile dependence.
Thus, as expected, and shown explicitly in Fig. 5, assumptions about the inner density profiles of UCMHs have a correspondingly mild impact on our constraints on the primordial power spectrum. We see that the constraints for steeper density profiles are slightly stronger at small scales, corresponding to the greater sensitivity to small masses in Fig. 4.
Figure 4: EDO bounds for different density profiles $\rho\propto r^{-\alpha}$ with fixed values of $R_{90}$ (blue) and the corresponding UCMH bounds (red). The solid lines are the same as in Fig. 1 in the main text.
Another aspect entering our limits is how we treat the UCMH mass. In the main text, we conservatively assume that $M_{\rm UCMH}$ is given by Eq. (1), with no subsequent growth due to accretion. While this behaviour corresponds to the finding in some simulations, other works have argued that $M_{\rm UCMH}$ should grow by a factor of $(1+z_{\rm eq})/(1+z_{c})$ during matter domination. We illustrate in Fig. 5 the impact of such a mass growth, concluding that the impact on the CMB accretion and Icarus limits on $\mathcal{P}_{\mathcal{R}}(k)$ is noticeable but still relatively small, while the impact on the wide binary evaporation limits is significant. In combination, this would imply that UCMH constraints outperform CMB spectral distortion constraints over the entire range of scales considered in our analysis. The reason why wide binary evaporation is particularly sensitive to the UCMH mass is simply that the constraint on $\mathcal{P}_{\mathcal{R}}(k)$ scales nonlinearly with the constraint on $f_{\rm DM}$, such that a fractional change in a less stringent constraint on $f_{\rm DM}$ affects $\mathcal{P}_{\mathcal{R}}(k)$ more strongly. We conclude our discussion of the potential impact of a UCMH mass growth by noting that a growth $\propto(1+z_{c})^{-1}$ can in any case not be the correct prescription for large values of $\mathcal{P}_{\mathcal{R}}(k)$: once there is a significant probability $\beta\gtrsim\mathcal{O}(10^{-3})$ for a region of a given size to collapse, cf. Eq. (3), there is no longer a sufficient amount of unbounded background matter in the surrounding regions to sustain such a mass growth (formally resulting in ‘overclosure constraints’ $\Omega_{\rm UCMH}>\Omega_{\rm DM}$ if applying the $\propto(1+z_{c})^{-1}$ scaling also in this regime).
Figure 5: Comparison of UCMH constraints for different inner density profiles and accretion mechanisms. Solid lines correspond to the limits presented in Fig. 2 in the main text. Dash-dotted lines show the same constraints assuming a density profile $\rho_{\rm UCMH}\propto r^{-\alpha}$ with $\alpha=9/4$ rather than $\alpha=3/2$. Dashed lines show the effect of a UCMH mass growth due to accretion, as explained in the text.
Appendix C Spectrum dependence of $\mu$-distortion and PTA constraints
In the main text we have pointed out that the CMB spectral distortion and PTA constraints typically probe a larger range of scales than UCMHs. In Fig. 6 we quantify this statement by showing the resulting constraints for a spectrum that is scale-invariant, $\mathcal{P}_{\mathcal{R}}(k)=const.$, over various different ranges of scales $k$ (while assuming $\mathcal{P}_{\mathcal{R}}(k)\simeq 0$ outside this range).
We stress that a fully scale-invariant power spectrum, $\mathcal{P}_{\mathcal{R}}(k)=const.$ for all $k$, is largely an academic exercise that we only include here for illustration: for normalizations larger than $\sim 10^{-9}$, both CMB spectral distortion and PTA constraints become irrelevant as such a choice is completely excluded by large-scale CMB temperature anisotropy observations. For direct comparison with the literature [28], we also include a (blue) line corresponding to the range of scales $0.15\lesssim k/k_{R}\lesssim 12.35$ that contribute $99$% to the mass variance $\sigma_{\chi,H}^{2}$ given in Eq. (4).
In Fig. 2 of the main text, we have presented results for a spectrum that is locally scale-invariant, implicitly assuming that $\mathcal{P}_{\mathcal{R}}(k)$ is non-vanishing only for a relatively small window in $k$ because the UCMH constraints are anyway only sensitive to scales $k\sim k_{R}$. Fig. 6 demonstrates that the UCMH constraints will clearly dominate over both CMB spectral distortion and PTA constraints even for much ‘broader’ power spectra, with $\mathcal{P}_{\mathcal{R}}(k)\simeq const.$ over a comparatively large range of scales.
Appendix D Constraints from Highly Magnified Stars
Observations of individual background stars near cluster caustics have opened a novel avenue for constraining compact dark matter. The smooth mass distribution of a galaxy cluster by itself produces a critical curve in the image plane and a corresponding fold caustic in the source plane whenever the projected surface density approaches the critical value. A secondary, smaller-scale lens, such as an intracluster (ICL) star or compact object?perturbs this macro caustic, creating a web of microcaustics around it. In favourable configurations, where a background star lies close to this combined macro?micro caustic, the magnification can reach extreme values ($\mu\sim 10^{3}$ or more), but the detailed light curve depends sensitively on whether the caustic remains smooth or is fragmented by compact lenses.
Figure 6: Comparison of the bounds from CMB $\mu$ distortions and PTAs for different integration ranges over a scale-invariant spectrum, as indicated. The blue line corresponds to a range $0.15\lesssim k/k_{R}\lesssim 12.35$, following Ref. [28].
As described in the main text, the first such constraint was derived from “Icarus” (MACS J1149 LS1 at $z=1.49$), whose magnification reached $\mu\sim 2000$  [71, 53, 54]. For compact lenses, three mass regimes are imporant: (i) for small lenses, the effective Einstein radius ($\bar{R}_{E}\propto\sqrt{M_{\rm EDO}}$) is smaller than the stellar radius, so the star smooths over microcaustics and no constraint arises; (ii) for intermediate masses, the magnification is reduced in a way that depends only on the total surface mass density in compact lenses, yielding a constraint on $f_{\rm DM}$ independent of $M$; (iii) for large $M$, the EDO itself acts as the main lens, and requiring a detectable caustic crossing implies $f_{\rm DM}\propto M^{1/3}$, with no constraint for $M\gtrsim 10^{4}M_{\odot}$.
A simplified analytic condition comes from the saturation of the critical curve by microcaustics. The effective Einstein radius of a compact lens is $\bar{R}_{E}\equiv\sqrt{\mu_{t}}\theta_{E}D_{L}$, such that the optical depth is
$\Theta=\frac{\Sigma}{M}\pi\bar{R}_{E}^{2},$ (5)
where $\Sigma$ is the surface mass density. A larger $\mu_{t}$ boosts $\bar{R}_{E}$, causing the effective lensing cross-section to increase and promoting microcaustic overlap even at small $f_{\rm DM}$. Setting $\Theta=1$ gives an upper limit on the macro-magnification for a given $f_{\rm DM}\lesssim f_{p}=\Sigma/\kappa\Sigma_{\rm crit}$ with $\Sigma_{\rm crit}\equiv M/\pi R_{E}^{2}=M/\pi\theta_{E}^{2}D_{L}^{2}$ with the Einstein angle $\theta_{E}$ and $D_{L}$ for Icarus given in the main text, and $\kappa\approx 0.83$ the best-fit fraction for the event: [^1] Here, $f_{p}$ quantifies the surface density in compact objects (which includes EDOs as well as ICL stars) normalised by the critical surface density. The constraint is conservative when interpreted as a bound on EDOs, since any known stellar contribution (e.g., ICL stars) would only enhance the microcaustic density and lead to a lower true value of $f_{\rm DM}$ for EDOs.
$\mu_{t,\rm max}=\frac{M}{\pi\Sigma R_{E}^{2}}\approx 1.2f_{\rm DM}^{-1}\,,$ (6)
where the last equality takes into account the distances to the lens and source based on the GLAFIC model used for MACS J1149 [72, 73]. To explain the Icarus observation with a $10M_{\odot}$ ICL star, $\mu_{t}>15.4$ is needed [53], which leads to $f_{\rm DM}\leq 7.8\times 10^{-2}$. Note that the great majority of ICL stars is $0.1-1\rm\,M_{\odot}$, so this constraint is conservative; if the ICL star is instead $0.1\rm\,M_{\odot}$ the constraint would be $f_{\rm DM}\leq 1.7\times 10^{-2}$, valid for $R_{90}\lesssim 4.3\times 10^{6}\bigl{(}M/M_{\odot}\bigr{)}^{1/2}R_{\odot}$. For EDOs larger than $10M_{\odot}$, brighter peak magnitudes are reached if the EDO plays the role of the lens.
Crucially, the derivation of the Icarus constraint in [53] on point-like lenses assumed point sources, an assumption which breaks down when $\bar{R}_{E}\lesssim R_{\star}$, as microcaustics are then averaged over by the source’s finite size. A more realistic treatment requires ray-tracing simulations with extended sources and finite-size lenses (e.g. [54]). In part of the regime, complementary constraints were derived based in the expected event rate [74].
Additional events promise to refine or strengthen these constraints. The “Godzilla” event in the Sunburst Arc ($z=2.37$) reached $\mu\sim 600$ with microcaustic peaks up to a few $10^{3}$ [75]; “Spock” transients in MACS J0416 showed $\mu\sim 400$ [76]; “Mothra” may exceed $\mu\sim 6000$ if its compact perturber is confirmed [77]. “Earendel” (at $z=6.2$) which has been found to have $\mu>4000$ from JWST imaging [78]. [^2] However, lens models in Ref. [79] suggest a lower value of $\mu\approx 43$-$67$. This discrepancy highlights the sensitivity of such constraints to lens modeling systematics and underscores the need for continued development of cluster mass reconstructions. These increasingly numerous caustic-crossing stars may offer a pathway to probe smaller compact objects than previously accessible. Since $\bar{R}_{E}\propto\sqrt{\mu_{t}}$, higher-magnification systems enhance sensitivity to lower-mass, as well as more extended, lenses. Future constraints will likely benefit from improved ray-tracing and lens modeling, as well as a statistical sample of such events across multiple clusters.